1. Below are the annualized yields on zero - coupon U.S. Treasuries. Use these data to answer the following questions.

Maturity 1-yr 2-yr 3-yr 5-yr 7-yr 10-yr 20-yr 30-yr
Rate 0.79% 1.15% 1.40% 1.81% 2.14% 2.34% 2.68% 2.96%

a. What is the seven-year forward holding rate of return between years 3 and 10 (r3,10)?

b. What is the annualized rate of return between years 7 and 30(r7,30​)?
Answer #1

a. 7 year forward rate = (1.023410/1.02143)1/7 - 1 = 2.43%

b. rate of return between years 7 and 30 = (1.029630/1.02147)1/23 - 1 = 3.21%